40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 14 13:29:51 2018, gautam, Update, General, PSL shutter closed, all optics misaligned 
    Reply  Thu Jun 14 19:30:12 2018, gautam, Update, General, IFO alignment restored 
       Reply  Thu Jun 14 22:45:05 2018, johannes, Update, General, AUX beam SRC alignment 
          Reply  Sat Jun 16 00:26:48 2018, gautam, Update, General, PRC modescan attempt 
             Reply  Sat Jun 16 01:25:29 2018, Koji, Update, General, PRC modescan attempt 
                Reply  Sat Jun 16 20:57:59 2018, Jon, Update, General, PRC modescan attempt 40M_SRMI_AM_annotated.pdf40M_DRMI_AM.pdf
Message ID: 13968     Entry time: Thu Jun 14 22:45:05 2018     In reply to: 13967     Reply to this: 13974
Author: johannes 
Type: Update 
Category: General 
Subject: AUX beam SRC alignment 

[Jon, Gautam, Johannes]

Jon spent some time trying to align the AUX beam to the SRC today, I got to the game kind of late so maybe others can add more detail.

The AUX beam that is reflected by the SRM looks terribly misshapen - it is quite elongated in vertical direction. Unfortunately I didn't snap a picture of it - anybody? It seemed at first as if this could be clipping - but after confirming the alignment of the AUX beam with the PSL output beam with aligned SRM, a slow dither of the SRM just moved the ugly pattern on the AS camera with no change to its shape - so clipping is unlikely. I'm now thinking that this is just the output beam of the fiber coupler after propagating ~15 meters to the SRM and back - even though this aspheric lens triplet coupler is supposed to be super-duper. I found that if I loosen the fiber slightly and pull it back just a bit at least the spot on the AS camera becomes nice and round - so maybe the fiber just doesn't sit well in this collimator? Not sure why that would be. I checked the fiber tip with the microscope, and while there was some gunk present, the central region and the core were clear (still cleaned using the fiber cleaning kit, which got rid of the debris). Either way, before switching to a different collimator I think we should give the Guoy phase measurement a shot - after all there was plenty of RF signal present on both AS110 and the PDA10CF placed at the YEND.

Looking for rogue beams on the AS table, I started placing some beam dumps. There was one particularly strong source of stray beams - a lens that was labeled with KPX094AR.33_F100. It became apparent after alignment efforts to the IFO had moved the AUX beam signifcantly off-center on this lens. According to the label it should have an AR coating for 1064nm, however judging by the amount of reflected light, it was certainly NOT AR-coated for 1064nm. I replaced it with a bi-convex f=100mm lens with confirmed AR-behavior.

The AUX laser is currently shuttered.


Per our Wednesday meeting, some items to work on are

  • Align the zero-order AUX beam into a second collimator on the PSL table, so we can switch the fiber output and look for RF signals at the offset-phaselock frequency without the additional frequency shift from the AOM. This will simpligy the mode spectroscopy scheme significantly
  • Abandon the R10/T90 beamsplitters in favor of R90/T10 beamsplitters. We'll swap the large mirror in front of the AS camera with an R90/T10 BS, and follow it up with a second R90/T10 BS that sends the AUX beam to the IFO. This way we'll have identical power levels on AS110 and AS55, and still 90% of the current AUX light going into the IFO, but without strong secondary beams from R10/T90 optics.
ELOG V3.1.3-